OpenStudy (anonymous):

A certain population of fruit flies which demonstrates exponential growth is 200 flies after 10 days. If the initial population was 15 flies, what will be the population after 3 weeks.

OpenStudy (anonymous):

Just apply the exponential population growth formula A = Pe^(rt) First apply it for the first 10 days to get the value of 'r'. Then apply it for 3 weeks (t = 3*7 = 21) to get the answer. Can you do it from this info?